# Low-Resolution Self-Attention for Semantic Segmentation

### v2

@@ 4.5 Advanced LRFormer with Query-based Decoders
We conduct the experiments in ADE20K dataset, following the same experimental setup. Results are shown in Tab. XII. LRFormer+ demonstrated superior performance outshining recent query-based frameworks such as Mask2Former and Mask DINO. We can observe that Mask DINO is 0.2% better than Mask2Former with additional 42G FLOPs. SeMask is a more efficient architecture, which surpasses Mask2Former by 0.5% with 4G more FLOPs. When comparing the upgraded LRFormer+ with a Mask2Former model utilizing the Swin-B backbone, LRFormer+ achieved 1.9% higher mIoU than Mask2Former, despite LRFormer+ having slightly more parameters but significantly 31G lower FLOPs, indicating a more efficient architecture. Furthermore, the implementation of Mask2Former with the innovative P2T backbone showcased enhanced capabilities, with the P2T-L variant reaching a mIoU of 52.5%, 1.2% better than Swin-S version of Mask2Former with similar FLOPs. Nonetheless, LRFormer+ still outperformed this configuration. For example, LRFormer-B+ is 1.2% further better than the enhanced P2T-L+ version.
For a more intuitive analysis, we visualized the accuracy-FLOPs comparisons of Tab. III and Tab. XII in Fig. 2. From the curve and the data points of this figure, LRFormer series achieve higher accuracy with fewer FLOPs compared to all other models, like Mask2Former [39], Mask DINO [42], and P2T [20].
@@ 5 Conclusion
Acknowledgements. This work is funded by NSFC (No. 62225604, 62176130), the Fundamental Research Funds for the Central Universities (Nankai University, 070-63233089), and A*STAR Career Development Fund (No. C233312006). Computation was partially supported by the Supercomputing Center of Nankai University. The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg).

### v3

@@ 4.5 Advanced LRFormer with Query-based Decoders
We conduct the experiments in the ADE20K dataset, following the same experimental setup. Results are shown in Tab. XIII. LRFormer+ demonstrated superior performance outshining recent query-based frameworks such as Mask2Former and Mask DINO. We can observe that Mask DINO is 0.2% better than Mask2Former with additional 42G FLOPs. SeMask is a more efficient architecture, which surpasses Mask2Former by 0.5% with 4G more FLOPs. When comparing the upgraded LRFormer+ with a Mask2Former model utilizing the Swin-B backbone, LRFormer+ achieved 1.9% higher mIoU than Mask2Former, despite LRFormer+ having slightly more parameters but significantly 31G lower FLOPs, indicating a more efficient architecture. Furthermore, the implementation of Mask2Former with the innovative P2T backbone showcased enhanced capabilities, with the P2T-L variant reaching a mIoU of 52.5%, 1.2% better than Swin-S version of Mask2Former with similar FLOPs. Nonetheless, LRFormer+ still outperformed this configuration. For example, LRFormer-B+ is 1.2% further better than the enhanced P2T-L+ version.
For a more intuitive analysis, we visualized the accuracy-FLOPs comparisons of Tab. III and Tab. XIII in Fig. 2. From the curve and the data points of this figure, LRFormer series achieve higher accuracy with fewer FLOPs compared to all other models, like Mask2Former [42], Mask DINO [45], and P2T [20].
TABLE XIV: Performance comparison of different backbones on the vision-language model LISA [91] for referring segmentation.
  Backbone gIoU (%) cIoU (%)
  ViT-L [15] 36.9 41.1
Swin-L [17] 38.1 43.1
LRFormer-XL (Ours) 40.9 45.7
4.6 Application to Vision-Language Models
While semantic segmentation remains fundamental in computer vision, the community has also shown growing interest in reasoning segmentation tasks. These new tasks integrate visual perception with language understanding capabilities [91, 92, 93, 94]. These tasks, exemplified by representative works like LISA [91], leverage large vision-language models such as CLIP [92] and LLaVA [93] to segment objects based on textual descriptions. To demonstrate the versatility of our proposed LRFormer beyond semantic segmentation, we conduct experiments on referring segmentation to verify whether our backbone can enhance the performance of vision-language models.
Experimental setup.  For our evaluation, we use LISA [91] with LLaVA-7B-v1 [93] as the baseline. We only replace the vision backbone of LISA [91] with three different options: ViT-L [15], Swin-L [17], and our LRFormer-XL, while keeping all other components consistent. Each backbone is pretrained on the COCO dataset [95] to ensure a fair comparison, and we maintain the same architecture for other parts of the vision branch. We use the official strategy [91] to train each method and test on the ReasonSeg validation set [91] for referring segmentation, which allows segmenting specific objects in images based on language prompts. Following previous works [96, 97, 91], we use gIoU and cIoU as evaluation metrics. More details of these metrics can refer to [91].
Results.  Tab. XIV shows the performance comparison of different backbones. Our LRFormer-XL backbone significantly outperforms both ViT-L [15] and Swin-L [17] across both metrics. Specifically, LRFormer-XL achieves 40.9% gIoU and 45.7% cIoU, surpassing the ViT-L [15] backbone by 4.0% and 4.6% in gIoU and cIoU, and the Swin-L [17] backbone by 2.8% and 2.6%, respectively. These results validate that our LRFormer effectively captures global context while preserving fine-grained details necessary for reasoning tasks. The consistent performance improvements across both conventional semantic segmentation and referring segmentation highlight the versatility and potential of our architecture for various advanced vision-language applications.
@@ 5 Conclusion
Acknowledgements. This work is funded by NSFC (No. 62225604, 62176130), the Fundamental Research Funds for the Central Universities (Nankai University, 070-63233089), and A*STAR Career Development Fund (No. C233312006), and National Research Foundation Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2023-007). Computation was partially supported by the Supercomputing Center of Nankai University. The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg).